# PII-Scope: A Benchmark for Training Data PII Leakage Assessment in LLMs

### v1

@@ 9 Research Directions
Why do PII Attacks Succeed? Numerous studies have examined the internal workings of LLMs from a safety perspective (Chen et al., 2024; Bereska & Gavves, 2024; Arditi et al., 2024). Few recent works have shifted the focus toward privacy concerns, identifying neurons responsible for data leakage (Wu et al., 2023), using activation steering techniques (Wu et al., 2024b), or exploring unlearning processes (Jang et al., 2022). A key limitation of these approaches is their reliance on simple zero-shot template attacks for evaluation (Huang et al., 2022a), raising concerns about the robustness of these interpretability-based mitigations. For example,  (Patil et al., 2023) shows that LLM unlearning does not fully erase private data, which can still be retrieved by probing internal layers (Patil et al., 2023). Furthermore, a recent work (Łucki et al., 2024) reveals that unlearning techniques (Li et al., 2024a) are prone to obfuscation, and a simple few-shot finetuning can restore unsafe capabilities. Therefore, a thorough analysis of privacy assessments against strong adversaries and an understanding of the underlying factors behind successful attacks is crucial.
@@ 10 Summary and Conclusion
In this work, we introduce PII-Scope, an empirical benchmark for assessing PII leakage from LLMs in different treat settings. We first evaluated the robustness of each PII attack method with respect to its internal hyperparameters. Our analysis uncovered key findings: hard-prompt attacks are highly sensitive to prompt structure and context, while soft-prompt attacks are influenced by prompt initialization and the number of training epochs. Furthermore, we demonstrated that PII attacks in a single-query setting significantly underestimate the extent of PII leakage. We show that attackers can exploit various combinations within these methods to launch multi-query attacks, and can dynamically adapt their strategies in continual settings, and achieve up to a 5.4x boost in extraction rates with modest query budgets.
Additionally, we compared the extraction rates of finetuned model to pretrained model, empirically demonstrating the significantly elevated privacy risks in finetuned settings. We achieved extraction rates exceeding 60% on the finetuned model with fewer than 500 queries. Overall, we hope that our work provides a fair and realistic benchmark for evaluating PII leakage, offering insights into how attackers can enhance extraction rates, and emphasizing the need for more robust defenses.
@@ 11 Limitations
Our work has several limitations. Firstly, recent LLMs such as LLama (Touvron et al., 2023) and Gemma (Team et al., 2023) do not disclose the sources of their pretraining datasets. Due to the lack of publicly available PII benchmark datasets that were part of these pretraining datasets, our analysis is confined to examining a single type of PII—email—found in the Enron email dataset (Shetty & Adibi, 2004). Additionally, our evaluations are limited to base LLMs and do not extend to instruction-tuned counterparts, i.e., aligned LLMs, which may exhibit different behaviors in response to PII extraction prompts. Specifically, in the aligned LLM setting, the focus shifts towards jailbreaking aligned models back to their base configurations, where the attacks discussed in this paper could be applied to extract PII from the jailbroken models. In the future, we plan to conduct an empirical evaluation of PII jailbreaking techniques on aligned LLMs (Touvron et al., 2023; Team et al., 2023).
@@ A.1 Reproducibility
Dataset. We provide the details of $M=64$ data subjects in $\mathcal{D}_{adv}$ in Figures 21 and  22, and the details of 308 data subjects in $\mathcal{D}_{eval}$ in Figures 29 and 30. Additionally, we conducted experiments with synthetic data subjects in $\mathcal{D}_{\text{adv}}^{s}$, where only the name part is anonymized (see Figures 23 and  24). In Figures 25 and  26, both the name and domain parts are anonymized.
@@
For the base configuration in SPT which we mentioned in § 5.5, we initialize the soft prompt embeddings with the embeddings of the task-aware string ‘‘Extract the email address associated with the given name’’ and set the number of soft-prompt embeddings $L$ to $50$. We train the soft prompt embeddings for $20$ epochs and report the best performance across all epochs. The training is conducted on the data subjects in the Adversary set $\mathcal{D}_{\text{adv}}$, containing $M=64$name, email} PII pairs ie., $\{s_{j}^{*},p_{j}^{*}\}_{j=1}^{M}$.
@@
Hyperparameters for Finetuning. We finetuned GPTJ-6B (Wang & Komatsuzaki, 2021) for two epochs with a batch size of 8. We used the AdamW optimizer (Loshchilov, 2017) with a learning rate of 0.0005 and a weight decay of 0.01. The original Enron email dataset (Shetty & Adibi, 2004), containing about 530K email bodies, was chunked into segments of 256 tokens. We then randomly selected 80% of the chunked data for finetuning.

### v2

@@ 13 Research Directions
Why do PII Attacks Succeed? Numerous studies have examined the internal workings of LLMs from a safety perspective (Chen et al., 2024; Bereska & Gavves, 2024; Arditi et al., 2024). Few recent works have shifted the focus toward privacy concerns, identifying neurons responsible for data leakage (Wu et al., 2023), using activation steering techniques (Wu et al., 2024b), or exploring unlearning processes (Jang et al., 2022). A key limitation of these approaches is their reliance on simple zero-shot template attacks for evaluation (Huang et al., 2022a), raising concerns about the robustness of these interpretability-based mitigations. For example,  (Patil et al., 2023) shows that LLM unlearning does not fully erase private data, which can still be retrieved by probing internal layers (Patil et al., 2023). Furthermore, a recent work (Łucki et al., 2024) reveals that unlearning techniques (Li et al., 2024a) are prone to obfuscation, and a simple few-shot finetuning can restore unsafe capabilities. Therefore, a thorough analysis of privacy assessments against strong adversaries and an understanding of the underlying factors behind successful attacks is crucial. Moreover, as demonstrated in our experiments, various factors—such as demonstrations in ICL (see Figure 7(b)), prepended prefixes (see Figure 7(c)) in the PII-Compass attack, and prompts constructed with synthetic data (see Figure 11(a))—influence the extraction rates. These insights will help pave the way for future mechanistic interpretations, providing a deeper understanding of the underlying factors driving this behavior.
@@ 14 Summary and Conclusion
In this work, we conducted an empirical benchmarking to assessing PII leakage from LLMs in different treat settings. We first evaluated the robustness of each PII attack method with respect to its internal hyperparameters. Our analysis uncovered key findings: hard-prompt attacks are highly sensitive to prompt structure and context, while soft-prompt attacks are influenced by prompt initialization and the number of training epochs. Furthermore, we demonstrated that PII attacks in a single-query setting significantly underestimate the extent of PII leakage. We show that attackers can exploit various combinations within these methods to launch multi-query attacks, and can dynamically adapt their strategies in continual settings, and achieve up to a five-fold boost in extraction rates for email PII with modest query budgets.
Additionally, we compared the extraction rates of finetuned models to pretrained models, empirically demonstrating the significantly elevated privacy risks in finetuned settings. Moreover, we have provided evidence of the underestimation of privacy leakage across different model families, including GPT-J, Pythia, and LLaMa models, as well as across two PIIs: email and phone number. Overall, we hope that our work provides a fair and realistic benchmark for evaluating PII leakage, offering insights into how attackers can enhance extraction rates, and emphasizing the need for more robust defenses.
@@ 15 Limitations
While our work provides a comprehensive evaluation of PII attacks, several limitations must be acknowledged.
  1. 1.
    Limited PII Evaluations: Recent open-source LLMs, such as LLaMa (Touvron et al., 2023), Phi Abdin et al. (2024), and Gemma (Team et al., 2023), do not disclose the sources of their pretraining datasets. As a result, the lack of publicly available PIIs from these datasets limits our analysis to just two types of PII: email and phone number, both found in the Enron email subset (Shetty & Adibi, 2004) of PILE. We have limited our evaluations to two models, GPT-J-6B Wang & Komatsuzaki (2021) and Pythia-6.9 (Biderman et al., 2023).
    Apart from email and phone number PII, which are part of the Enron email dataset, we were unable to identify other PIIs (e.g., social security numbers, passport numbers) within the pretrained datasets of popular models. This further limits the thorough assessment of privacy leakage across different entities in these models’ pretrained data.
  2. 2.
    Non-Instruction Models: Our evaluations are limited to base LLMs and do not extend to instruction-tuned models (i.e., aligned LLMs), which may exhibit different behaviors in response to PII extraction prompts. Specifically, in the aligned LLM setting, the focus shifts to jailbreaking the models back to their base configurations using prompt-engineering techniques.
    In the future, we plan to empirically evaluate PII jailbreaking techniques, such as AutoDAN (Liu et al., 2023) and PAIR (Chao et al., 2023), on aligned LLMs (Touvron et al., 2023; Team et al., 2023) to extract PIIs. Additionally, we aim to extend PII attacks to other entities in LLMs fine-tuned with synthetic PII datasets.
16 Broader Impact Concerns
Although our work demonstrates evidence of increased extraction rates that could potentially aid attackers, we hope it raises awareness about the significant underestimation of privacy leakage in LLMs, particularly in the context of single-query attacks. By benchmarking against attacks with repeated queries and in continual settings, we provide a valuable evaluation framework for LLM providers to audit their models in-house and gain a better understanding of the extent of privacy leakage. Our empirical benchmarking of PII attacks, especially through repeated querying, highlights a critical vulnerability in current systems that requires further attention and improvement.
Furthermore, our work contributes to a deeper understanding of the memorization of PII tokens in LLMs, which requires further study to understand the mechanistic interpretation of why a given prompt (e.g., ICL prompt, PII-Compass prompt from different subjects, or soft prompt) aids in PII extraction. The implications of this research extend beyond academia and could encourage practitioners to more thoroughly audit the privacy risks associated with LLMs. By fostering this deeper understanding, we aim to inspire further advancements in privacy-preserving techniques to defend against the strongest adversaries and contribute to the overall enhancement of LLM privacy.
@@ A.1 Reproducibility
Similarly, we extract phone numbers from the generated string using the below regex expression.
import re
[…]
Email PII Dataset. We provide the details of $M=64$ data subjects in $\mathcal{D}_{adv}$ in Figures 24 and  25, and the details of 308 data subjects in $\mathcal{D}_{eval}$ in Figures 32 and 33. Additionally, we conducted experiments with synthetic data subjects in $\mathcal{D}_{\text{adv}}^{s}$, where only the name part is anonymized (see Figures 26 and  27). In Figures 28 and  29, both the name and domain parts are anonymized.
Phone Number PII Dataset. We provide the details of $M=64$ data subjects in $\mathcal{D}_{adv}$ in Figures 14(a) and  14(a), and the details of 436 data subjects in $\mathcal{D}_{eval}$ in Figures 14(a) and 14(a).
@@
For the base configuration in SPT which we mentioned in § 6.5, we initialize the soft prompt embeddings with the embeddings of the task-aware string ‘‘Extract the <PII> associated with the given name’’ and set the number of soft-prompt embeddings $L$ to $50$. We train the soft prompt embeddings for $20$ epochs and report the best performance across all epochs. The training is conducted on the data subjects in the Adversary set $\mathcal{D}_{\text{adv}}$, containing $M=64$name, PII entity} pairs ie., $\{s_{j}^{*},p_{j}^{*}\}_{j=1}^{M}$.
@@
Hyperparameters for Finetuning. We finetuned GPTJ-6B (Wang & Komatsuzaki, 2021) and Pythia 6.9B (Biderman et al., 2023) for two epochs with a batch size of 8. We used the AdamW optimizer (Loshchilov, 2017) with a learning rate of 0.0005 and a weight decay of 0.01. The original Enron email dataset (Shetty & Adibi, 2004), containing about 530K email bodies, was chunked into segments of 256 tokens. We then randomly selected 80% of the chunked data for finetuning. For Llama7B (Touvron et al., 2023) experiments, we used the publicly available finetuned models which is finetuned for 10 epochs, provided by authors of LLM-PBE (Li et al., 2024b).